OpenStudy (anonymous):

A particle is moving along the curve y=3sqrt(2x+5). As the particle passes through the point (2,9) , its x-coordinate increases at a rate of units per second. Find the rate of change of the distance from the particle to the origin at this instant.

OpenStudy (anonymous):

ur missing dx/dt in your problem....

OpenStudy (anonymous):

ohh its increasing at a rate of 2 units per second

OpenStudy (anonymous):

let d = distance from the point to the origin... so: \(\large d=\sqrt{x^2+y^2} \) or \(\large d^2=x^2+y^2 \) since \(\large y=3\sqrt{2x+5} \), the equation becomes: \(\large d^2=x^2+9(2x+5)^2 \) can you differentiate wrt time here?

OpenStudy (anonymous):

im not sure a little confused

OpenStudy (anonymous):

2d=2x+9(4x^2+10x+25) ?

OpenStudy (anonymous):

2d=2x+9(8x+10) ?

OpenStudy (anonymous):

maybe simplifying it a bit will help..... \(\large d^2=x^2+9(2x+5)^2 \) \(\large d^2=x^2+9(4x^2+20x+25) \) \(\large d^2=x^2+36x^2+180x+225 \) \(\large d^2=37x^2+180x+225 \) so differentiating wrt time: \(\large [d^2]'=[37x^2+180x+225]' \) \(\large 2dd'=74xx'+180x' \) is this ok ??? lemme know if you need explanation of any step...

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

since you're given x', and x, you'll need to find d.... do you know how to do that?

OpenStudy (anonymous):

sub in for x and x prime?

OpenStudy (anonymous):

not yet... find d first....

OpenStudy (anonymous):

when x=2, y=9 (the point is given) find the distance from this point to the origin.

OpenStudy (anonymous):

how do we do that?

OpenStudy (anonymous):

use the formula above... \(\large d=\sqrt{x^2+y^2} \)

OpenStudy (anonymous):

\(\large d=\sqrt{(\color {red}2)^2+(\color {red}9)^2} \) = ???

OpenStudy (anonymous):

once you have d, you can substitute that value in along with x' into the equation: \(\large 2dd'=74xx'+180x' \rightarrow d'=\frac{74xx'+180x'}{2d} \)

OpenStudy (anonymous):

so now you substitute?

OpenStudy (anonymous):

d=sqrt 85

OpenStudy (anonymous):

ok... good... so now just use the formula for d' (the rate at which the distance to the origin is changing) by replacing x=2, x'=2, d=sqrt(85)...

OpenStudy (anonymous):

= 74(2)(2)+180*2/2sqrt)85) ?

OpenStudy (anonymous):

= [74(2)(2)+180*2]/[2sqrt(85)] if that's what you meant, then yes....